OpenStudy (anonymous):

60% of a number is 90. What is 50% of that same number? A. 30 B. 45 C. 75 D. 150

OpenStudy (kobeni-chan):

First you want to find where the 60% came from. Divide 90 by 0.60 for the first number.

OpenStudy (kobeni-chan):

Once you have that number, you can either divide by 2 or multiply by 0.50 (0.5) to find 50% of the same number :)

OpenStudy (anonymous):

yes i know what 50 50 means hal

OpenStudy (kobeni-chan):

did you find what the original number was?

OpenStudy (anonymous):

150

OpenStudy (kobeni-chan):

Good :) so the answer is half (50%) of that.

OpenStudy (anonymous):

75

OpenStudy (kobeni-chan):

right! c:

OpenStudy (anonymous):

thanks

OpenStudy (kobeni-chan):

no problem

OpenStudy (xapproachesinfinity):

60% percent of a number is 90 could be interpreted in algebra as follows: \[\large \frac{60}{100}x=90\]

OpenStudy (xapproachesinfinity):

60% is the same as 60/100 of means product is means equal

OpenStudy (xapproachesinfinity):

solve for x

OpenStudy (xapproachesinfinity):

so if solve for x: \[\large x=\frac{90.100}{60}=\frac{900}{6}=150\] now you need to find 50% of that

OpenStudy (xapproachesinfinity):

this a systematic way of doing it!